find the length of the hypoteneuse of a right triangle if its two legs are 6m and 8m?
Answer by nerdybill(7384)

find the length of the hypoteneuse of a right triangle if its two legs are 6m and 8m?
.
Apply the Pythagorean theorem:
Let h = hypotenuse
then
h^2 = 6^2 + 8^2
h^2 = 36 + 64
h^2 = 100
h = 10m